OpenStudy (loser66):

Generalize help P is polynomials, prove T is a linear transformation T(p(x))=p(x^2) I can do in particular one, but not know how to make it in general please, guide me

OpenStudy (loser66):

p(x1) =x+2 --> T(p(x1)= x^2 +2 p(x2) = x^3+2x -->T(p(x2) = x^6+2x^2 T(p(x1)+p(x2))= T(px1)+T(px2) adding a proof of scalar multiplication, I can get T is a linear transformation But I need a general form and don't know how to do

OpenStudy (loser66):

@wio

OpenStudy (anonymous):

Well, \(p(x)\) is allegedly a polynomial. \[ p(x) = \sum_{k}a_xx^k \]

OpenStudy (loser66):

yes

OpenStudy (loser66):

a_k x^k, typo, right?

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Here is one:\[ T(cp(x)) = cp(x^2) = c[p(x^2)] = cT(p(x)) \]

OpenStudy (loser66):

I mean p(x) = sum a_k x^k

OpenStudy (anonymous):

\[ T(p_1(x) + p_2(x)) = p_1(x^2)+p_2(x^2) = T(p_1(x)) + T(p_2(x)) \]

OpenStudy (loser66):

how can we go from left to middle one?

OpenStudy (anonymous):

Let \(p_3(x) = p_1(x)+p_2(x)\) Then \(p_3(x^2) = p_1(x^2)+p_2(x^2)\)

OpenStudy (loser66):

oh yeaaaah. Thanks a ton